# Towards Scalable Exact Machine Unlearning Using Parameter-Efficient Fine-Tuning

### v2

@@ Abstract
Machine unlearning is the process of efficiently removing the influence of a training data instance from a trained machine learning model without retraining it from scratch. A popular subclass of unlearning approaches is exact machine unlearning, which focuses on techniques that explicitly guarantee the removal of the influence of a data instance from a model. Exact unlearning approaches use a machine learning model in which individual components are trained on disjoint subsets of the data. During deletion, exact unlearning approaches only retrain the affected components rather than the entire model. While existing approaches reduce retraining costs, it can still be expensive for an organization to retrain a model component as it requires halting a system in production, which leads to service failure and adversely impacts customers. To address these challenges, we introduce an exact unlearning framework – Sequence-aware Sharded Sliced Training (S3T), which is designed to enhance the deletion capabilities of an exact unlearning system while minimizing the impact on model’s performance. At the core of S3T, we utilize a lightweight parameter-efficient fine-tuning approach that enables parameter isolation by sequentially training layers with disjoint data slices. This enables efficient unlearning by simply deactivating the layers affected by data deletion. Furthermore, to reduce the retraining cost and improve model performance, we train the model on multiple data sequences, which allows S3T to handle an increased number of deletion requests. Both theoretically and empirically, we demonstrate that S3T attains superior deletion capabilities and enhanced performance compared to baselines across a wide range of settings.
@@ 2 Background
Figure 1: Schematic diagram of the Sharded, Isolated, Sliced, and Aggregated training (SISA) (Bourtoule et al., 2021) framework. An ensemble of models is individually trained on disjoint shards. (Left) Each shard is further divided into slices. (Right) Each model is sequentially trained on the slices and checkpoints are stored. After deletion, retraining resumes from the best available checkpoint.
Exact Machine Unlearning. These techniques focus on developing a modular machine learning system, where individual components are trained using disjoint subsets of the data. Such a system offers the advantage that when a deletion request is received for an input instance, we only need to retrain the affected component rather than the entire model. However, these systems require modifying the original training process of the model. The seminal work for such a modular unlearning system is Sharded, Isolated, Sliced, and Aggregated training (SISA) (Bourtoule et al., 2021). SISA uses an ensemble of models each trained on a disjoint shard of the dataset as shown in Figure 1 (left). To further reduce retraining costs, each shard is divided into slices, and the models are incrementally trained on these slices, with their checkpoints stored sequentially (shown in Figure 1 (right)). If a deletion request affects a data instance within the 4th slice of a shard, we must retrieve the checkpoint after the 3rd slice and retrain using the remaining data within that shard. Several approaches focus on improving the components within SISA in application-specific settings like enhancing the dataset partitioning mechanism (Aldaghri et al., 2021; Yan et al., 2022), the retraining efficiency using light-weight adapters (Kumar et al., 2023; Dukler et al., 2023), or extending the modular approach for vision tasks by using compartmentalized diffusion models (Golatkar et al., 2023). However, these approaches are orthogonal to our work since they do not fundamentally modify the functioning of SISA. These modifications can be easily incorporated within our proposed framework.
@@ 3.3 Training under Budget Constraints
The training strategy discussed in the previous section requires training the same model on $L!$ slice sequences, which is expensive in practice. In this section, we discuss strategies to select sequences under a budget, $B$ (maximum number of sequences that can trained). First, we show that there exists an optimal subset of $B$ slice sequences and randomly selecting $B$ sequences may not be effective. To illustrate this idea, we use a permutation tree (Bhattacharya, 1994), where all possible permutation sequences are embedded into a tree structure.
In Figure 3 (left), we show an example of a permutation tree with $L=3$ slices, paths from the root to the leaves correspond to unique permutation sequences $(S_{1},S_{2},S_{3})$, $(S_{1},S_{3},S_{2})$, and so on. From the previous section, we know that the topmost slice is the most sensitive because if a deletion request affects the topmost slice the entire model needs to be retrained (shown in Figure 2 (right)). To address this and reduce the retraining cost, we should ensure we train models on sequences with different top slices. Building on this intuition, in the general setting we should train the model on diverse permutation sequences. Two sequences are considered diverse if no element appears at the same position, e.g., $(S_{1},S_{2},S_{3})$ and $(S_{2},S_{3},S_{1})$. An example is illustrated in Figure 3 (left), where a diverse set of 3 sequences is marked in green (where no identical slices occupy the same position). Selecting diverse permutations is challenging as relying solely on random sampling may not always yield the best results. Moreover, in certain scenarios, it is possible to have prior knowledge about the deletion probabilities of data slices; for example, younger users might be more likely to request data deletion than older users. Therefore, we present two strategies for selecting diverse permutation sequences for a budget $B$, depending on whether or not prior deletion probabilities are available.
@@ 4.2 Deletion Performance
Next, we extensively evaluate the impact of increasing the training budget $B$ on the deletion rate (with $m=5$ shards & $L=64$ slices). In Figure 6 (right), we observe that there is a steady growth in the deletion rate with an increasing budget. The growth is slightly higher in the initial stages when the budget is low and slows down gradually. This experiment provides empirical evidence to our theoretical result in Lemma 2, which claims that the performance improves with an increased budget.
@@ 4.3 Analysis
Figure 8: We evaluate the performance of iterative cyclic rotation and bipartite matching-based selection (BMS). (Left) We observe that cyclic rotation selection consistently outperforms random sampling for all budgets, $1\leq B\leq 120$ (with fixed $L=5$). (Center) We evaluate the average edit distance of the sequences generated by BMS and observe that it achieves the optimal edit distance ($L$). (Right) We also observe that sequences from BMS achieves higher scores than random sampling.
@@
Sequence Selection. We evaluate the quality of the sequences generated by the iterative cyclic rotation and BMS algorithm (Section 3.3). Ideally, we want the selected sequences to be diverse and should have a high edit distance between sequences. Therefore, we report the average edit distance within a selected subset, $O$: $\mathbb{E}_{o\in O}[d_{\mathrm{edit}}(o,o^{\prime})]$. First, we consider the setting where the deletion prior is uniform and compare cyclic rotation with random sampling. In Figure 8 (left), we report the average edit distance with varying budget ($B$) while the slice count ($L$) is fixed. We observe that cyclic rotation produces significantly better sequences than random sampling. Second, we consider slices associated with a deletion prior. We present the results by averaging over 10 deletion priors sampled from a Dirichlet distribution. In Figure 8 (center & right), we observe that BMS consistently outperforms random sampling both in terms of diversity (avg. edit distance) and chosen sequence scores ($\mathrm{score}[\mathcal{S},t]$).
Deletion Time. Unlearning in S3T is cost-effective; it primarily involves selecting the best checkpoint and swapping with the current one. The main cost is incurred when a large number of deletion requests necessitate re-training the system from scratch. In Figure 9, we report the total deletion time as the number of deletion requests increases, which includes the checkpoint swapping and re-training time. In this experiment, we compare the deletion time of S3T with SISA and full re-training on CIFAR10 dataset. Full re-training retrains the model after each deletion request. For SISA and S3T, the step jumps indicate that the system requires retraining. Overall, S3T achieves the lowest deletion time. S3T reduces the total deletion time (after 1000 requests) by 2.8x compared to SISA and 25x compared to full retraining. This experiment shows the efficacy of S3T in reducing the overall deletion time.
@@ Acknowledgments
The authors thank Anneliese Brei, Anvesh Rao, Archiki Prasad, Haoyuan Li, Nicholas Monath, Rajeev Ambati, Rakesh Menon, Shounak Chattopadhyay, and Vinayshekhar Bannihatti Kumar for helpful discussions and feedback on the draft.
@@ A.1 Proof of Lemma 1
In S3T, since the sequences are selected to be diverse (Figure 4), the topmost slice in each sequence is different. Therefore, we have $B^{\prime}=\min(B,L)$ different slices at the topmost position. This implies that for S3T to encounter service failure, a total of $mB^{\prime}$ slices must be affected by deletion requests, where $m$ is the number of shards. Considering deletion requests affect all slices uniformly we need to compute the expected time till all slices are affected. This setup is similar to the coupon collector problem (Blom et al., 1994).
@@
where $\mathbb{E}[t_{i}]=1/p_{i}$, where $p_{i}$ is the probability that the $i$-th slice is affected after $(i-1)$ slices are deleted. Let $N=|\mathcal{D}|$ the dataset size and $r$ be the number of deletion requests seen so far. The expression of $p_{i}$ is shown below:
@@
The second part of the lemma is about SISA. For SISA to experience failure, only the first slice of each shard (total of $m$ slices) needs to be affected by deletion. In this case, we can write:
@@ C.3 Additional Experiments
Figure 15: We compare the scores (Eq. 24) of the generated sequences by BMS and sorted cyclic rotation. We observe BMS consistently outperforms cyclic rotation.
BMS vs. Cyclic rotation with Deletion prior. In this setting, we compare the BMS selection algorithm with a variant of cyclic rotation when the prior deletion probabilities are available. We perform cyclic rotation by first sorting the slices based on their deletion probabilities. For example, we sort the sequence with deletion probabilities: ($S_{1}$: 0.5, $S_{2}$: 0.4, $S_{3}$: 0.1) as: ($S_{3},S_{2},S_{1}$). Then, for a budget $B=3$, the stored sequences are: ($S_{3},S_{2},S_{1}$), ($S_{1},S_{3},S_{2}$), ($S_{2},S_{1},S_{3}$). In this variant, the slices most likely to be deleted are not at the top of any sequences. We follow the experimental setup in Section 4.3 and sample deletion priors using a Dirichlet distribution (over 10 runs). In Figure 15, we report the total sequence scores (Eq. 24) obtained by BMS and sorted cyclic rotation for a budget, $B=L$. We observe that BMS consistently outperforms sorted cyclic rotation for all budgets. Please note that the average edit distance achieved by both methods are the same as cyclic rotations are guaranteed to produce the maximum diversity for budgets: $B\leq L$.
@@
Appendix D Broader Impact

### v3

@@ Abstract
Machine unlearning is the process of efficiently removing the influence of a training data instance from a trained machine learning model without retraining it from scratch. A popular subclass of unlearning approaches is exact machine unlearning, which focuses on techniques that explicitly guarantee the removal of the influence of a data instance from a model. Exact unlearning approaches use a machine learning model in which individual components are trained on disjoint subsets of the data. During deletion, exact unlearning approaches only retrain the affected components rather than the entire model. While existing approaches reduce retraining costs, it can still be expensive for an organization to retrain a model component as it requires halting a system in production, which leads to service failure and adversely impacts customers. To address these challenges, we introduce an exact unlearning framework – Sequence-aware Sharded Sliced Training (S3T), which is designed to enhance the deletion capabilities of an exact unlearning system while minimizing the impact on model’s performance. At the core of S3T, we utilize a lightweight parameter-efficient fine-tuning approach that enables parameter isolation by sequentially training layers with disjoint data slices. This enables efficient unlearning by simply deactivating the layers affected by data deletion. Furthermore, to reduce the retraining cost and improve model performance, we train the model on multiple data sequences, which allows S3T to handle an increased number of deletion requests. Both theoretically and empirically, we demonstrate that S3T attains superior deletion capabilities and enhanced performance compared to baselines across a wide range of settings. [^1] https://github.com/brcsomnath/S3T
@@ 2 Background
Figure 1: Schematic diagram of the Sharded, Isolated, Sliced, and Aggregated training (SISA) (Bourtoule et al., 2021) framework. An ensemble of models is individually trained on disjoint shards ($\mathcal{D}_{i}$’s). (Left) Each shard is further divided into slices. (Right) Each model is sequentially trained on the slices and checkpoints are stored. After deletion, retraining resumes from the best available checkpoint.
Exact Machine Unlearning. These techniques focus on developing a modular machine learning system, where individual components are trained using disjoint subsets of the data. Such a system offers the advantage that when a deletion request is received for an input instance, we only need to retrain the affected component rather than the entire model. Exact unlearning systems can guarantee unlearning and privacy as none of its components is affected by the deleted instance. However, these systems require modifying the original training process of the model. The seminal work for such a modular unlearning system is Sharded, Isolated, Sliced, and Aggregated training (SISA) (Bourtoule et al., 2021). SISA uses an ensemble of models each trained on a disjoint shard of the dataset as shown in Figure 1 (left). To further reduce retraining costs, each shard is divided into slices, and the models are incrementally trained on these slices, with their checkpoints stored sequentially (shown in Figure 1 (right)). If a deletion request affects a data instance within the 4th slice of a shard, we must retrieve the checkpoint after the 3rd slice and retrain using the remaining data within that shard. Several approaches focus on improving the components within SISA in application-specific settings like enhancing the dataset partitioning mechanism (Aldaghri et al., 2021; Yan et al., 2022), the retraining efficiency using light-weight adapters (Kumar et al., 2023; Dukler et al., 2023), or extending the modular approach for vision tasks by using compartmentalized diffusion models (Golatkar et al., 2023). However, these approaches are orthogonal to our work since they do not fundamentally modify the functioning of SISA. These modifications can be easily incorporated within our framework.
@@ 3.3 Training under Budget Constraints
The training strategy discussed in the previous section requires training the same model on $L!$ slice sequences, which is expensive in practice. In this section, we discuss strategies to select sequences under a budget, $B$ (maximum number of sequences that can be trained). First, we show that an optimal subset of $B$ slice sequences exists, and randomly selecting $B$ sequences may not be effective. To illustrate this idea, we use a permutation tree (Bhattacharya, 1994), where all possible permutation sequences are embedded into a tree structure.
In Figure 3 (left), we show an example of a permutation tree with $L=3$ slices, paths from the root to the leaves correspond to unique permutation sequences $(S_{1},S_{2},S_{3})$, $(S_{1},S_{3},S_{2})$, and so on. From the previous section, we know that the topmost slice is the most sensitive because if a deletion request affects the topmost slice, the entire model needs to be retrained (shown in Figure 2 (right)). To address this and reduce the retraining cost, we should ensure we train models on sequences with different top slices. Building on this intuition, in the general setting we should train the model on diverse permutation sequences. Two sequences are considered diverse if no element appears at the same position, e.g., $(S_{1},S_{2},S_{3})$ and $(S_{2},S_{3},S_{1})$. An example is illustrated in Figure 3 (left), where a diverse set of 3 sequences is marked in green (where no identical slices occupy the same position). Selecting diverse permutations is challenging as relying solely on random sampling may not always yield the best results. Moreover, in certain scenarios, it is possible to have prior knowledge about the deletion probabilities of data slices; for example, younger users might be more likely to request data deletion than older users. Therefore, we present two strategies for selecting diverse permutation sequences for a budget $B$, depending on whether or not prior deletion probabilities are available.
@@ 4.2 Deletion Performance
Next, we extensively evaluate the impact of increasing the training budget $B$ on the deletion rate (with $m=5$ shards & $L=64$ slices). In Figure 6 (right), we observe that there is a steady growth in the deletion rate with an increasing budget. The growth is slightly higher in the initial stages when the budget is low and slows down gradually. This experiment helps answer (RQ2) and provides empirical evidence to our theoretical result in Lemma 2, which claims that the performance improves with an increased budget.
@@ 4.3 Analysis
Figure 8: We evaluate the performance of iterative cyclic rotation and bipartite matching-based selection (BMS). (Left) We observe that cyclic rotation selection consistently outperforms random sampling for all budgets, $1\leq B\leq 120$ (with fixed $L=5$). (Center) We evaluate the average edit distance of the sequences generated by BMS and observe that it achieves the optimal edit distance ($L$). (Right) We also observe that sequences from BMS achieve higher scores than random sampling.
@@
Sequence Selection. We evaluate the quality of the sequences generated by the iterative cyclic rotation and BMS algorithm (Section 3.3). Ideally, we want the selected sequences to be diverse and should have a high edit distance between sequences. Therefore, we report the average edit distance within a selected subset, $O$: $\mathbb{E}_{o\in O}[d_{\mathrm{edit}}(o,o^{\prime})]$. First, we consider the setting where the deletion prior is uniform and compare cyclic rotation with random sampling. In Figure 8 (left), we report the average edit distance with varying budget ($B$) while the slice count ($L$) is fixed. We observe that cyclic rotation produces significantly better sequences than random sampling. Second, we consider slices associated with a deletion prior. We present the results by averaging over 10 deletion priors sampled from a Dirichlet distribution. In Figure 8 (center & right), we observe that BMS consistently outperforms random sampling both in terms of diversity (avg. edit distance) and chosen sequence scores ($\mathrm{score}[\mathcal{S},t]$). These experiments help answer (RQ3), showcasing the effectiveness of the proposed sequence selection algorithms in practice.
Deletion Time. Unlearning in S3T is cost-effective; it primarily involves selecting the best checkpoint and swapping it with the current one. The main cost is incurred when a large number of deletion requests necessitate re-training from scratch. In Figure 9, we report the total deletion time (in hours) as the number of deletion requests increases, which includes the checkpoint swapping and re-training time. In this experiment, we compare the deletion time of S3T with SISA and full re-training on the CIFAR10 dataset. Full re-training retrains the model after each deletion request. For SISA and S3T, the step jumps indicate that the system requires retraining. Overall, S3T achieves the lowest deletion time. S3T reduces the total deletion time (after 1000 requests) by 2.8x compared to SISA and 25x compared to full retraining. This experiment shows the efficacy of S3T in reducing the deletion time.
@@ 5 Conclusion
Reproducibility Statement
We have submitted the implementation of S3T in the supplementary materials. We have extensively discussed the details of our experimental setup, datasets, baselines, and hyperparameters in Section 4 and Appendix C. We have also provided the details of our training procedure and sequence selection algorithms in Appendix B. We will make our implementation public once the paper is published.
Ethics Statement
In this paper, we introduce an exact unlearning framework, S3T, to enable the efficient unlearning of training data from large-scale models. S3T has been developed to enable users to remove the impact of private and harmful data from ML models. However, it is crucial to assess whether the deletion requests are initiated by legitimate users. An adversarial user could exploit this feature to degrade the model’s performance. In our work, we use publicly available datasets and open-sourced models.
@@ Acknowledgments
The authors thank Anneliese Brei, Anvesh Rao, Archiki Prasad, Haoyuan Li, Nicholas Monath, Rajeev Ambati, Rakesh Menon, Shounak Chattopadhyay, and Vinayshekhar Bannihatti Kumar for helpful discussions and feedback on the draft. This work was supported in part by NSF grant DRL-2112635.
@@ A.1 Proof of Lemma 1
In S3T, since the sequences are selected to be diverse (Figure 4), the topmost slice in each sequence is different. Therefore, we have $B^{\prime}=\min(B,L)$ different slices at the topmost position. This implies that for S3T to encounter service failure, a total of $mB^{\prime}$ slices must be affected by deletion requests, where $m$ is the number of shards. Considering deletion requests affect all slices uniformly, we need to compute the expected time till all slices are affected. This setup is similar to the coupon collector problem (Blom et al., 1994).
@@
where $\mathbb{E}[t_{i}]=1/p_{i}$, where $p_{i}$ is the probability that the $i$-th slice is affected after $(i-1)$ slices are deleted. Let $N=|\mathcal{D}|$ be the dataset size and $r$ be the number of deletion requests seen so far. The expression of $p_{i}$ is shown below:
@@
The second part of the lemma is about SISA. For SISA to fail, only the first slice of each shard (total of $m$ slices) needs to be affected by deletion. In this case, we can write:
@@ C.3 Additional Experiments
Figure 15: (Left) We compare the scores (Eq. 24) of the generated sequences by BMS and sorted cyclic rotation. We observe BMS consistently outperforms cyclic rotation. (Center) We show the variation in deletion rate with increasing budget $B$ under non-uniform deletion prior over the slices. (Right) We report the performance of SISA and protoSISA with varying weights.
BMS vs. Cyclic rotation with Deletion prior. In this setting, we compare the BMS selection algorithm with a variant of cyclic rotation when the prior deletion probabilities are available. We perform cyclic rotation by first sorting the slices based on their deletion probabilities. For example, we sort the sequence with deletion probabilities: ($S_{1}$: 0.5, $S_{2}$: 0.4, $S_{3}$: 0.1) as: ($S_{3},S_{2},S_{1}$). Then, for a budget $B=3$, the stored sequences are: ($S_{3},S_{2},S_{1}$), ($S_{1},S_{3},S_{2}$), ($S_{2},S_{1},S_{3}$). In this variant, the slices most likely to be deleted are not at the top of any sequences. We follow the experimental setup in Section 4.3 and sample deletion priors using a Dirichlet distribution (over 10 runs). In Figure 15 (left), we report the total sequence scores (Eq. 24) obtained by BMS and sorted cyclic rotation for a budget, $B=L$. We observe that BMS consistently outperforms sorted cyclic rotation for all budgets. Please note that the average edit distance achieved by both methods is the same as cyclic rotations are guaranteed to produce the maximum diversity for budgets: $B\leq L$.
Deletion with Non-uniform Prior. In this setting, we evaluate the impact of increasing the budget $B$ when the deletion prior is non-uniform. We use $m=5$ shards and $L=32$ slices. For each shard, we sample a deletion prior over the slices distributions from a Dirichlet distribution. Using these priors, we select sequences to train the model using the BMS algorithm. Since the BMS algorithm always generates $L$ sequences, we greedily select the top $B$ sequences based on the sequence score defined in Eq. 24. In Figure 15 (center), we report the deletion rate of S3T under varying budget sizes, $B$, when the deletion requests are sampled from the prior. We observe that when we have access to the deletion prior, S3T can handle significantly more deletion requests (compared to the uniform prior setting in Figure 6 (right)). Overall, the deletion rate shows a gradual increase as the budget rises; however, it remains notably high even with a low budget.
Figure 16: We report the model performance using an ensemble approach where individual models are trained on different slice sequences. We observe an initial improvement in performance, which eventually plateaus.
Baseline Ablations. In this setting, we investigate the performance of ProtoSISA under different settings. ProtoSISA uses the same setup as SISA along with a prototypical classifier. During inference, we use a weighted combination of the logits of SISA and the prototypical classifier. In this experiment, we vary the weight assigned to the SISA component and report the results in Figure 15 (right). We observe that ProtoSISA achieves slightly better performance with lower weights when the number of deletion requests is high.
Ensemble of Slice Sequences. In this experiment, we evaluate the impact of using an ensemble of models trained on different slice sequences. These slice sequences include data within the same shard. We use a ViT-base model on the CIFAR100 dataset ($m=5,L=4$). In Figure 16, we observe that increasing the number of models initially enhances performance; however, this improvement plateaus as the model count increases.
Appendix D Limitations & Future Directions
In this paper, we present a novel exact unlearning framework, S3T. S3T improves the deletion rate of existing unlearning systems at the cost of additional offline training. The training process can be time-consuming if we are fine-tuning larger models and have a higher performance requirement (thereby high budget $B$). However, this is an inherent tradeoff between offline training and losing revenue due to re-training costs. The developer should adjust their budget according to the tradeoff in their specific application.
The primary goal of an exact unlearning system is to maintain a finely modular structure, enabling efficient retraining of individual components upon data deletion. Therefore, it is important to have sample efficient learning methods so that components achieve high performance while using less data. To further improve S3T, this would involve research to develop more expressive parameter-efficient fine-tuning approaches that can perform well when trained with small amounts of data.
Appendix E Broader Impact